OpenStudy (anonymous):

how do i find the slope of y=-2/3x-2

OpenStudy (precal):

number in front of x is your slope

OpenStudy (precal):

y=mx+b m is your slope

OpenStudy (anonymous):

like i no the equation is y=mx +b right?

OpenStudy (precal):

y=mx+b is just what is known as slope and y intercept form

OpenStudy (anonymous):

so would it be y = -2/3x + -2 ??

OpenStudy (anonymous):

well how do i identify the slope and y-intercept form

OpenStudy (precal):